# Modeling Story Expectations to Understand Engagement: A Generative Framework Using LLMs

### v2

@@ 1 Introduction
We apply our method to a dataset of 30,258 book chapters we collected from Wattpad, a large online media platform that allows writers to publish their stories and readers to consume stories. We find that our method complements existing feature engineering techniques in improving model explanatory power in the following sense: if $f$ is a feature engineering technique that converts an unstructured story into a low-dimensional feature such that $f(ExistingStory)$ is useful for explaining engagement, then this same feature engineering technique can be applied to the imagined story continuations to generate a set of complementary features (i.e., expectations, $\mathbb{E}[f(ImaginedStories)]$, uncertainty, $Var[f(ImaginedStories)]$, and surprise, $Surprise[f(ImaginedStories)]$) that can further help explain engagement. Intuitively, if customers are likely to care about certain dimensions of the story that they have read and experienced, they are also likely to care about the expectations, uncertainty, and surprise along such dimensions. We quantify this complementarity by calculating the relative improvement in model performance from adding features based on imagined stories compared to adding features based on actual stories. We document that our framework amplifies the usefulness of the existing feature engineering techniques by around 31%.
@@ 3.1 Story Input
Our starting point is a piece of narrative content, such as the first few chapters of a book. One challenge in using LLMs is there is a limit on the size of the context window for the input into the model. For example, GPT-3.5-turbo can at most accommodate 16,385 tokens, which translates to roughly 12,300 words. [^2] https://platform.openai.com/docs/models/gpt-3-5-turbo, https://platform.openai.com/tokenizer While we can easily feed in the text for one chapter, the context window size becomes binding after just a few chapters. To overcome this issue, we use as input a summary of chapters $1$ to $t-1$ and the full text of chapter $t$ to generate the predicted stories.
@@
    Prompt for Chapter 1: “You are an average book reader. Here is the first chapter of a book. Provide an extensive summary of the chapter. Focus on the characters, their actions and emotions, and events in a coherent and consistent way without making assumptions. Don’t start with sentences mentioning the chapter or the book — such as ‘In this chapter’; get right into the summary.”
@@
    Prompt for Chapters 2+: “You are an average book reader. Here are the summary of the previous chapters of a book you have read, and the entire text of a new chapter. Provide an extensive summary of the entire book based on the previous chapters and the new chapter. Focus on the characters, their actions and emotions, and events in a coherent and consistent way without making assumptions. Don’t start with sentences mentioning the chapter or the book — such as ‘In this chapter’; get right into the summary.”
@@ 4 Data
To demonstrate our proposed methodology, we collect book text with chapter-by-chapter engagement. We collect data from Wattpad, an online media platform that allows users to read and write stories. In October 2024, Wattpad had over 90 million readers and writers. For readers, the vast majority of stories are free to read but some stories require a premium membership or payment to access the later chapters. The major advantage of using Wattpad is that content is posted chapter by chapter and we can observe the read count, vote count, and comment count for each chapter and therefore calculate the continue-to-read, vote-to-read, and comment-to-read rates.
We collected a corpus of publicly accessible free books from Wattpad by focusing on the genres listed on Wattpad’s homepage (e.g., action, adventure, romance). Appendix A provides additional details about the keywords used to collect the books. For each book, we collect its title and description, when the book was created, the language the book is written in, whether the content is for mature audiences, and writer-provided book tags (e.g., “school”, “drama”, “friendship”). For each chapter, we collect its text, title, date it was written, and its comment count, vote count, and read count.
We use GPT-4o-mini to summarize the chapters and GPT-3.5-turbo to generate the imagined story continuations. [^4] Specifically, we use GPT-4o-mini-2024-07-18 and GPT-3.5-turbo-0125. We use GPT-4o mini for text summarization because of its good summarization capabilities, its lower cost relative to GPT-3.5-turbo, and because we are not concerned about data leakage for this step. However, for the imagination generation, to alleviate the concern about data leakage (i.e., the engagement metrics and actual story continuations being part of the training data), we only include books that were published after the cutoff date of the training data for GPT-3.5-turbo (September 2021) so all book chapters are published January 2022 and onwards. This ensures the content was not used to train the GPT model we use to generate the imagined stories. Since our proposed story imagination method relies critically on the quality of the input text, we take several steps to clean the collected stories, which we detail in Appendix LABEL:appendix:wattpad_cleaning. After cleaning, we are left with 30,258 chapters across 1,735 books.
@@ 4.1 Summary Statistics
Table 1 provides summary statistics of our dataset of book chapters. Compared to more traditional books, which have on average 3,000 to 4,000 words per chapter, the Wattpad chapters are shorter with an average of 1,827 words per chapter. The engagement with these chapters is fairly high not only in terms of continuation rates but also in terms of comment and vote rates. Readers can comment throughout the text and comments are consolidated at the end of the chapter text. Readers can also vote to show their support for a writer and can only vote once per chapter. Our outcome measures of interest are:
@@ 5 Empirical Application with Books Data
We apply our proposed framework to the collected Wattpad data.
@@ 5.2.2 Regression Results
These findings suggest actionable implications for content creators and digital platforms aiming to optimize user engagement. Notably, the distinction between types of engagement – continued consumption versus active engagement like commenting and voting – points to an opportunity for differentiated engagement strategies. For example, if comments help to generate a sense of community and this is important for new customers, then a platform may want to recommend early on more positive and exciting content that generates expectations that the story will become more negative. On Wattpad, authors have the option to directly address their readers in their writing and can strategically place nudges to encourage readers to vote for or comment on their content based on the actual and anticipated emotion.
@@ 7 Conclusion
At the core of our methodology is a process that transforms narrative content into features representing audience expectations, uncertainty, and surprise. This approach bridges the gap between qualitative narrative analysis and quantitative modeling. To demonstrate its effectiveness, we applied our method to a dataset of over 30,000 book chapters from Wattpad. Our method complements existing feature engineering techniques by providing a framework to extend their application to reader beliefs, amplifying their marginal explanatory value by 31%. Through careful analysis of the regression results, we can uncover key narrative elements that correspond to audience engagement. These findings contribute to a deeper understanding of the relationship between narrative structure and audience response and generate hypotheses that can be further explored. In conclusion, our framework advances our ability to model audience expectations for narrative content. By quantifying consumers’ forward-looking beliefs, we provide a valuable tool for marketers, content creators, and researchers to better understand audience behavior.
@@ Appendix A Appendix: Data Cleaning
Our Wattpad dataset is collected based on specific search keywords related to the genres listed on Wattpad’s homepage. The genres listed are: Action, Adventure, ChickLit, Classics, Fanfiction, Fantasy, General Fiction, Historical Fiction, Horror, Humor, Mystery, Non-Fiction, Paranormal, Poetry, Random, Romance, Science Fiction, Short Story, Spiritual, Teen Fiction, Thriller, Vampire, Werewolf.
We preprocess the dataset at three levels—text, chapter, and book—to focus on content that best reflects narrative structure and reader engagement. This multi-stage cleaning ensures high-quality inputs for downstream analysis and generative modeling.
At the text level, we remove content elements that are clearly outside the core narrative. This includes author commentary, notes directed at readers, and other meta-text that may appear at the beginning or end of chapters. We also remove any formatting tags (e.g., HTML) and other non-linguistic artifacts that may interfere with language model processing.
Chapters are then filtered to retain only those likely to contain meaningful narrative content. We begin by identifying and excluding chapters that are unlikely to be part of the story—such as playlists, author updates, and character lists—using a combination of heuristics (e.g., titles, word count thresholds) and a language model classifier that assigns a likelihood score based on the text and metadata. Chapters with very low predicted narrative probability are excluded. To comply with content policies and ensure model compatibility, we further exclude chapters predicted to contain explicit material. These predictions are made using a language model that estimates the likelihood of such content. To avoid discontinuities in behavioral signals, we also exclude the chapter that immediately precedes flagged ones. We additionally remove chapters that appear to have been rewritten during the data collection window, which can reset engagement counters and distort metrics such as continuation rates. Chapters written shortly before the data collection date are excluded as well, due to limited exposure time and low engagement signal. We also omit final chapters (which lack follow-up continuation metrics) and chapters with zero read counts, which can skew rate-based calculations.
After chapter-level filtering, we apply further criteria at the book level. First, we restrict the dataset to English-language content, estimating language consistency by measuring the proportion of English words in each chapter and computing a book-level average. Books and chapters that fall below language thresholds are removed. We also exclude books tagged with mature content labels, as well as books with a high proportion of chapters predicted to contain sensitive content. This avoids gaps in narrative flow and ensures compatibility with content generation policies. Books with extremely short average chapter lengths (e.g., meme collections or short-form entries) and excessively long chapters (which may reflect full novels pasted into a single entry) are also removed. To maintain consistency in modeling and avoid overrepresentation, we drop books with more than 50 chapters. Finally, we exclude books whose first chapter shows minimal engagement, as this limits our ability to observe reader behavior.

### v3

@@ 1 Introduction
We apply our method to a dataset of over 30,000 book chapters we collected from a large online media platform that allows writers to publish their stories and readers to consume stories. We find that our method complements existing feature engineering techniques in improving model explanatory power in the following sense: if $f$ is a feature engineering technique that converts an unstructured story into a low-dimensional feature such that $f(ExistingStory)$ is useful for explaining engagement, then this same feature engineering technique can be applied to the imagined story continuations to generate a set of complementary features (i.e., expectations, $\mathbb{E}[f(ImaginedStories)]$, uncertainty, $Var[f(ImaginedStories)]$, and surprise, $Surprise[f(ImaginedStories)]$) that can further help explain engagement. Intuitively, if customers are likely to care about certain dimensions of the story that they have read and experienced, they are also likely to care about the expectations, uncertainty, and surprise along such dimensions. We quantify this complementarity by calculating the relative improvement in model performance from adding features based on imagined stories compared to adding features based on actual stories. We document that our framework amplifies the usefulness of the existing feature engineering techniques by around 31%.
@@ 3.1 Story Input
Our starting point is a piece of narrative content, such as the first few chapters of a book. One challenge in using LLMs is there is a limit on the size of the context window for the input into the model. For example, GPT-3.5-turbo can at most accommodate 16,385 tokens, which translates to roughly 12,300 words. [^2] https://platform.openai.com/docs/models/gpt-3-5-turbo, https://platform.openai.com/tokenizer While we can easily feed in the text for one chapter, the context window size becomes binding after just a few chapters. To overcome this issue, we use as input a summary of chapters 111 to $t-1$ and the full text of chapter $t$ to generate the predicted stories.
@@
    System Prompt: “You are an average book reader.”
@@
    User Prompt:
    1. (a)
      Prompt for Chapter 1: “Here is the first chapter of a book. Provide an extensive summary of the chapter. Focus on the characters, their actions and emotions, and events in a coherent and consistent way without making assumptions. Don’t start with sentences mentioning the chapter or the book — such as ‘In this chapter’; get right into the summary.”
    2. (b)
      Prompt for Chapters 2+: “Here are the summary of the previous chapters of a book you have read, and the entire text of a new chapter. Provide an extensive summary of the entire book based on the previous chapters and the new chapter. Focus on the characters, their actions and emotions, and events in a coherent and consistent way without making assumptions. Don’t start with sentences mentioning the chapter or the book — such as ‘In this chapter’; get right into the summary.”
@@ 3.2 Imagination Generation
    System Prompt: “You are an average book reader.”
  2. 2.
    User Prompt:
    1. (a)
@@ 4 Data
To demonstrate our proposed methodology, we collect book text with chapter-by-chapter engagement. We collect data from an online media platform that allows users to read and write stories. The content is posted chapter by chapter and we can observe the read count, vote count, and comment count for each chapter and therefore calculate the continue-to-read, vote-to-read, and comment-to-read rates.
For each book, we collect its title and description, when the book was created, the language the book is written in, whether the content is for mature audiences, and writer-provided book tags (e.g., “school”, “drama”, “friendship”). For each chapter, we collect its text, title, date it was written, and its comment count, vote count, and read count.
We use GPT-4o-mini to summarize the chapters and GPT-3.5-turbo to generate the imagined story continuations. [^4] Specifically, we use GPT-4o-mini-2024-07-18 and GPT-3.5-turbo-0125. We use GPT-4o mini for text summarization because of its good summarization capabilities, its lower cost relative to GPT-3.5-turbo, and because we are not concerned about data leakage for this step. However, for the imagination generation, to alleviate the concern about data leakage (i.e., the engagement metrics and actual story continuations being part of the training data), we only include books that were published after the cutoff date of the training data for GPT-3.5-turbo (September 2021) so all book chapters are published January 2022 and onwards. This ensures the content was not used to train the GPT model we use to generate the imagined stories. Since our proposed story imagination method relies critically on the quality of the input text, we take several steps to clean the stories, which we detail in Appendix A. After cleaning, we are left with 30,258 chapters across 1,735 books.
@@ 4.1 Summary Statistics
Table 1 provides summary statistics of our dataset of book chapters. Compared to more traditional books, which have on average 3,000 to 4,000 words per chapter, the chapters in our data are shorter with an average of 1,827 words per chapter. The engagement with these chapters is fairly high not only in terms of continuation rates but also in terms of comment and vote rates. Readers can comment throughout the text and comments are consolidated at the end of the chapter text. Readers can also vote to show their support for a writer and can only vote once per chapter. Our outcome measures of interest are:
@@ 5 Empirical Application with Books Data
We apply our proposed framework to the collected data.
@@ 5.2.2 Regression Results
These findings suggest actionable implications for content creators and digital platforms aiming to optimize user engagement. Notably, the distinction between types of engagement – continued consumption versus active engagement like commenting and voting – points to an opportunity for differentiated engagement strategies. For example, if comments help to generate a sense of community and this is important for new customers, then a platform may want to recommend early on more positive and exciting content that generates expectations that the story will become more negative.
@@ 7 Conclusion
At the core of our methodology is a process that transforms narrative content into features representing audience expectations, uncertainty, and surprise. This approach bridges the gap between qualitative narrative analysis and quantitative modeling. To demonstrate its effectiveness, we applied our method to a dataset of over 30,000 book chapters from a large online media platform. Our method complements existing feature engineering techniques by providing a framework to extend their application to reader beliefs, amplifying their marginal explanatory value by 31%. Through careful analysis of the regression results, we can uncover key narrative elements that correspond to audience engagement. These findings contribute to a deeper understanding of the relationship between narrative structure and audience response and generate hypotheses that can be further explored. In conclusion, our framework advances our ability to model audience expectations for narrative content. By quantifying consumers’ forward-looking beliefs, we provide a valuable tool for marketers, content creators, and researchers to better understand audience behavior.
@@ Appendix A Appendix: Data Cleaning
We clean the stories at the text-level, the chapter-level, and the book-level.
A.1 Cleaning the Text
We remove non-story related text. Specifically, we do the following:
  1. 1.
    Remove Special Notes: Sometimes writers directly address readers through notes at the beginning or end of the chapter. We remove them as they are not part of the narrative.
  2. 2.
    Remove HTML Tags and Attributes: We remove HTML tags and attributes since they are not part of the narrative.
A.2 Cleaning the Chapters
After cleaning the text, we remove chapters that are unrelated to the narrative (e.g., preludes, suggested music playlists), conflict with GPT’s text generation policies, or were rewritten, which restarts the read count and distorts the continue-to-read rate.
  1. 1.
    Remove Irrelevant Chapters: We remove chapters that are not part of the narrative (e.g., author’s notes, playlist suggestions, lists of characters) using several strategies. First, we remove chapters that are titled “introduction,” “prologue,” or “character” or contain the keywords “author” or “extra.” Second, we remove chapters that have a word count less than 2 standard deviations from the average word count of the book since these chapters are unlikely to be part of the narrative. Third, we prompt GPT to provide a score predicting whether the chapter is part of the story or not. In addition to the chapter text, we provide to GPT the chapter title and URL because they can also be informative. We remove chapters that return a probability score less than 0.25, a threshold based on manual inspection of the chapter text. We use the following prompt to determine whether a chapter is part of the story:
    “Please classify the following text and provide a probability score between 0 and 1. The score should be close to 1 if the text is a book chapter, and closer to 0 if the text is a prelude, author’s note, Spotify playlist, prologue, or anything else that is not a book chapter. Your response may consist exclusively of a number between 0 and 1. The chapter title is: ‘{title}’. The chapter url is: ‘{url}’.”
  2. 2.
    Remove Explicit Content: We remove chapters with explicit content since generating explicit content violates OpenAI’s usage policies. To identify explicit content, we prompt GPT to provide a score predicting whether the the chapter contains explicit content. We remove chapters that return a probability score greater than 0.9 as well as the preceding chapter since the continue-to-read rate will be affected. We use the following prompt to determine whether there is explicit content:
    “Evaluate the following text and assist me in determining if it features sexually explicit content. Your response may consist exclusively of a probability score from 0 to 1, where a score close to 1 indicates high confidence that the text contains sexually explicit material.”
  3. 3.
    Remove Rewritten Chapters: Sometimes authors rewrite chapters, which resets the read count for those chapters. We drop chapters that are rewritten as well as the preceding chapter since the continue-to-read rates will reflect the reset read count.
  4. 4.
    Remove Recently Written Chapters: The data was collected July 2024. We remove chapters created after May 2024 because new chapters are likely to provide noisy measures of engagement.
  5. 5.
    Remove Final Chapter: We remove the last chapter of each book since they do not have a continue-to-read rate. This also removes books with only one chapter.
  6. 6.
    Remove Chapters with Zero Read Count: We remove chapters that have a read count of zero because the continue-to-read rate will be infinity.
A.3 Cleaning the Books
Finally, the following steps outline which books are removed from the dataset:
  1. 1.
    Remove Non-English Books: We focus on books written in English because the features we derive from the author-written and GPT-imagined stories require English text. We calculate the percentage of words which are in English for each chapter. We then characterize each book by the average percentage of English words across all chapters. Books that have an average less than 73% are dropped from the dataset. In addition, we drop chapters that are less than 50% English.
  2. 2.
    Remove Mature Books: In addition to dropping explicit chapters, we drop books that are tagged as “mature” by the author and books that contain explicit content tags (e.g., “adultromance”). We also remove books that have many chapters flagged as explicit since this would lead to large gaps in the narrative. Specifically, we drop books that have more than 15% of its chapters removed due to being explicit content.
  3. 3.
    Remove Books Unlikely to Have a Narrative: Books with low average word counts tend to lack a narrative structure. For example, several books are just collections of memes. We remove books with an average word count per chapter of less than 400.
  4. 4.
    Remove Books with High Word Count: We remove books with an average word count per chapter greater than 10,000 because GPT is limited in the length of its input and because often these chapters are not chapters but instead entire books. For the same reason, we also remove chapters with more than 10,000 words.
  5. 5.
    Remove Books with More Than 50 Chapters: We remove books with a high number of chapters so that books with more chapters do not receive greater weight in our subsequent analysis.
  6. 6.
    Remove Books with First Chapter Read Count Less Than 100: We remove books with a low read count for the first chapter.